OpenStudy (anonymous):

if a and b are the roots of ax^2 + bx + c =0 then the value of a^3 + b^3/(a^2 + b^2)=?

mathslover (mathslover):

nice question wait

OpenStudy (anonymous):

just notice \[a^2+b^2=(a+b)^2-2ab \ \ \ and \\ a^3+b^3=(a+b)(a^2+b^2-ab)=((a+b)^2-3ab)\]

OpenStudy (ash2326):

nice hint @mukushla @Yahoo! do you know the relation between coefficients of a quadratic and its roots?

OpenStudy (anonymous):

a+b = -b/a

OpenStudy (anonymous):

ab = c/a

OpenStudy (ash2326):

Good :) Could you use the hints given by @mukushla to find your answer?

OpenStudy (anonymous):

got it thxxxxx

OpenStudy (ash2326):

@Yahoo! @mukushla deserves the medal, I don't ;)